# First Measurement of Charged Current Muon Neutrino-Induced $K^{+}$ Production on Argon using the MicroBooNE Detector

### v2

@@ Abstract
The MicroBooNE experiment is an 85 tonne active mass liquid argon time projection chamber neutrino detector exposed to the on-axis Booster Neutrino Beam (BNB) at Fermilab. One of MicroBooNE’s physics goals is the precise measurement of neutrino interactions on argon in the 1 GeV energy regime. Building on the capabilities of the MicroBooNE detector, this analysis identifies $K^{+}$ mesons, a key signature for the study of strange particle production in neutrino interactions. This measurement is furthermore valuable for background estimation for future nucleon decay searches and for improved reconstruction and particle identification capabilities in experiments such as the Deep Underground Neutrino Experiment (DUNE). In this letter, we present the first-ever measurement of a flux-integrated cross section for charged-current muon neutrino induced $K^{+}$ production on argon nuclei, determined to be 7.93 $\pm$ 3.27 (stat.) $\pm$ 2.92 (syst.) ×$\times~{}10^{-42}\;$ cm2/nucleon based on an analysis of 6.88×$\times 10^{20}$ protons on target.
@@
Upcoming large long-baseline neutrino oscillation experiments, such as the Deep Underground Neutrino Experiment (DUNE) Abi et al. (2020) and Hyper-Kamiokande Abe et al. (2018), have the sensitivity to search for proton decay in this range. Cherenkov detector experiments like Super-Kamiokande Abe et al. (2014, 2017, 2020) and Hyper-Kamiokande are particularly suited for $p\rightarrow e^{+}\pi^{0}$ decay searches. However, for water Cherenkov detectors, the $p\rightarrow\nu K^{+}$ decay channel needs to be indirectly detected since the momentum of the produced $K^{+}$ (339 MeV/c) is below the Cherenkov detection threshold in water (560 MeV/c). Furthermore, backgrounds from atmospheric neutrino interactions, such as the $\nu_{\mu}n\rightarrow\mu^{-}K^{+}n$ and $\nu p\rightarrow\nu K^{+}\Lambda$ channels, are challenging to distinguish in water Cherenkov detectors.
@@
[…]
As the first quantitative measurement of $K^{+}$ production on argon nuclei, this analysis will impact nuclear, neutrino, and proton decay research on two fronts. First, it provides background constraints for future nucleon decay experiments. Atmospheric neutrinos can also produce a $K^{+}$ along with other particles, mimicking the signature of $p\rightarrow\nu K^{+}$ decay. Thus, neutrino-induced $K^{+}$ production serves as a background for proton decay studies. Second, this measurement provides input to improve neutrino generator models. In the case of bound proton decay, the $K^{+}$ is generated within the nucleus and may interact with other nucleons as it escapes. These interactions can decrease the detected momentum of the $K^{+}$ in the detector and may also eject nucleons from the nucleus.
Charged-current neutrino-induced $K^{+}$ production occurs through three interaction modes: (1) single kaon production, where a CC interaction produces a single kaon with a production energy threshold of $E_{\nu_{\mu}}$geq 0.79 GeV, and strangeness is not conserved ($|\Delta S|=1$); (2) associated kaon production, where both a kaon and a hyperon are produced, with a threshold of $E_{\nu_{\mu}}$geq 1.2 GeV, and strangeness is conserved; and (3) coherent production, where only one kaon is produced in the final state, and the target nucleus remains intact after the neutrino interaction.
Figure 1 shows an event candidate of CC $\nu_{\mu}$ kaon production recorded in MicroBooNE data. The kaon decays at rest into a 388 MeV energy muon, which subsequently decays into a Michel electron.
@@
The MicroBooNE detector Fleming (2012); Acciarri et al. (2017) consists of an 85-tonne LArTPC and a photon detection system comprised of 32 photomultiplier tubes (PMTs). The detector identifies neutrino interactions by detecting ionization and scintillation light produced by charged particles moving through the detector. Ionization charge is recorded across three wire planes (two induction, one collection) oriented at different angles (with induction planes at ±$\pm 60^{\circ}$ relative to the vertical collection plane). This configuration allows the generation of high-resolution, three-dimensional images of neutrino interactions. Scintillation light collected by the PMTs provides precise timing information, essential for identifying neutrino interactions that are synchronized with the BNB, and for rejecting cosmic-ray backgrounds.
@@
In MicroBooNE, the LArSoft framework Snider and Petrillo (2017) provides tools for simulating the neutrino flux, neutrino interactions, particle propagation, and detector response including ionization and scintillation processes as well as reconstruction of neutrino events. The BNB neutrino flux through the MicroBooNE detector is simulated using the MiniBooNE flux model Aguilar-Arevalo et al. (2009) adapted for MicroBooNE’s location along the beamline. GENIE v3.0.6 (G18_10a_02_11a) Andreopoulos et al. (2010, 2021) is the event generator used to simulate neutrino-argon interactions, including both single and associated kaon production. In GENIE, associated kaons are produced via individual resonances simulated by the Rein-Sehgal model Rein and Sehgal (1981). Kaons produced via Deep Inelastic Scattering (DIS) are simulated by the Koba-Nielsen-Olesen (KNO) scaling parameterization Koba et al. (1972). For higher energies (hadronic invariant mass W $>$ 2.3 GeV), the AGKY model Yang et al. (2009) and PYTHIA6 Sjöstrand et al. (2006) are used. The hadronization models are tuned to match strange particle production ($\Lambda$ and $K^{0}_{S}$) observed in experiments such as the Big European Bubble Chamber (BEBC) Jones et al. (1993); Allasia et al. (1983); Willocq et al. (1992); Bosetti et al. (1982) and the Fermilab bubble chamber Baker et al. (1986); DeProspo et al. (1994). The single kaon production model Rafi Alam et al. (2010) is implemented in GENIE, generating events in the channels $\nu_{l}+n\rightarrow l^{-}+n+K^{+}$, $\nu_{l}+p\rightarrow l^{-}+p+K^{+}$, and $\nu_{l}+n\rightarrow l^{-}+p+K^{0}$. The region of validity assumed for the model is for $E_{\nu}\leq$ 2 GeV Rafi Alam et al. (2010); Alvarez-Ruso et al. (2013), while the GENIE authors note that the model can be applied to higher energies as well Alam et al. (2015).
@@
Particle trajectories and interactions within the detector are simulated using GEANT4 Agostinelli et al. (2003), which models particle propagation in the detector. Simulated neutrino interactions are overlayed with data events collected using an unbiased trigger that operates in anti-coincidence with the beam, enabling data-driven modeling of cosmic ray and detector noise. Both simulation and data reconstruction use the Pandora framework Acciarri et al. (2018), which uses a multi-algorithm approach to pattern recognition and event reconstruction in LArTPCs. Pandora creates clusters from wire hits, identifies the neutrino vertex, matches clusters across wire planes to form 3D objects, classifies reconstructed particles as tracks or showers, and establishes particle hierarchies.
@@
For the $K^{+}$ event selection, we implemented a boosted decision tree (BDT) classifier Hoecker et al. (2009). The BDT classifier inputs include reconstructed variables such as $\chi^{2}$ values which compare the expected $dE/dx$ for kaon ($\chi^{2}_{k}$), proton ($\chi^{2}_{p}$), pion ($\chi^{2}_{\pi}$), and muon ($\chi^{2}_{\mu}$) hypotheses with the observed data for each wire plane, the combined 3-wire-plane $\chi^{2}$, the daughter track length, and the log-likelihood ratio particle identification (LLR PID) Abratenko et al. (2021). We trained the BDT classifier to identify true $K^{+}$ that decay to $\mu^{+}\nu_{\mu}$ or $\pi^{+}\pi^{0}$, using Monte Carlo (MC) simulation samples of reconstructed neutrino interactions that meet the preselection criteria described above. To train the BDT classifier, a background sample corresponding to $3.67\times 10^{20}$ POT (equivalent to 10% of the MicroBooNE BNB simulation) was used. For the simulation of the signal, a dedicated high-statistics sample of single and associated $K^{+}$ production was used. We optimized the BDT classifier score selection criteria by maximizing the product of efficiency and purity. Figure 2 shows the BDT classifier score distribution, with the orange arrow at 0.41 representing the selected BDT score cut (BDT selection). Events with a score higher than the BDT selection are kept as $K^{+}$ candidates, and 10 such $K^{+}$ candidates are identified in the data. The total number of background events predicted from the overlay simulation and off-beam data is 2.21 events. The simulation predicts a selection efficiency of 3.95%, and an associated selection purity of 89%. The primary background category for $K^{+}$ candidates consists of protons generated at the neutrino vertex that interact with an argon nucleus, producing secondary particles at the proton endpoint that could mimic a $K^{+}$ decay product, consisting of 2.08 events.
@@
Figure 3: Reconstructed daughter track length distribution. The stacked histogram shows the MC simulation with its statistical uncertainty (gray band). The black dots represent the MicroBooNE data. The red triangles are the $K^{+}$ candidates selected in MicroBooNE data by the BDT.
Figure 3 shows the reconstructed daughter track length distribution of the $K^{+}$ candidates selected in MicroBooNE data, including the event shown in Fig. 1. The stacked histogram shows the expected daughter track length distribution from the MicroBooNE BNB simulation. Due to the two-body decay nature of the $K^{+}$ branching ratios analyzed in this letter, the resulting mono-energetic $\mu^{+}$ ($\pi^{+}$) provides a powerful discriminant, as it leads to a fixed track length that can be measured with high accuracy and is not present in background events. The expected track length for $\mu^{+}$ and $\pi^{+}$ accumulates around 53 cm and 30 cm, respectively. The $K^{+}$ candidates selected by the BDT selection in Fig. 3 (red triangles) exhibit a pile-up in the region compatible with $K^{+}$ decay.
To evaluate the accuracy of the background modeling, two sideband regions were defined using a 2-dimensional distribution of the BDT classifier score versus daughter track length. A far sideband region is defined to include tracks with low BDT classifier scores (BDT $<$ 0) and track lengths less than 40 cm. A near sideband region selects tracks with BDT classifier scores close to the signal region (BDT $<$ 0.41) and across all track length ranges, but excludes tracks in the far sideband region. The data/MC comparisons of the near and far sidebands show agreement within uncertainties for the reconstructed kinematic variables such as $K^{+}$ candidate track length, $\phi$ (azimuthal angle around the beam direction), and $\theta$ (angle of the track with respect to the beam direction). The agreement with the data suggests that the simulation is able to adequately model the background in the cross section extraction. More information about the sideband comparisons is provided in the Supplemental Material sup .
This analysis presents the measurement of the single-bin flux-integrated cross section for $\nu_{\mu}+\text{Ar}\rightarrow\mu^{-}+K^{+}+X$ interactions, where muon neutrinos interact with Ar and produce one $\mu^{-}$, one $K^{+}$ and $X$ (any number of other hadrons in the final state) on argon nuclei. This measurement uses an exposure of $6.88\times 10^{20}$ POT from the on-axis BNB neutrino beam. The cross section ($\sigma$) was found to be 7.93 $\pm$ 3.27 (stat.) $\pm$ 2.92 (syst.) $10^{-42}\text{cm}^{2}/{\rm nucleon}$, and it was calculated using the following expression:
@@
where $N$ represents the number of selected signal events (10), $B$ is the number of expected background events (2.21), $\varepsilon$ is the selection efficiency ($3.95\%$), $N_{\text{target}}$ is the number of target nucleons ($4.13\times 10^{31}$), and $\Phi_{\nu_{\mu}}$ stands for the integrated BNB $\nu_{\mu}$ flux ($6.02\times 10^{11}\nu/\text{cm}^{2}$). The selection efficiency accounts for all possible $K^{+}$ decays within the fiducial volume, while the BDT selection identifies $K^{+}$ candidates that could decay into either $\mu^{+}$ or $\pi^{+}$.
[…]
@@
where $\rho$ is the density of the material, $x$ is the track length, $\sigma^{\text{tot}}_{\text{geant}}$ and $\sigma^{\text{inel}}_{\text{geant}}$ are the $K^{+}$ total and inelastic GEANT4 cross sections, and $\sigma^{\text{tot}}_{\text{var}}$ and $\sigma^{\text{inel}}_{\text{var}}$ are the $K^{+}$ total and inelastic cross sections estimated by applying a variation of $\pm$40to the $\sigma^{\text{inel}}_{\text{geant}}$. The 40% variation covers the disagreement between data and GEANT4 in $K^{+}-\text{Ar}$ interaction measurements reported by the LArIAT experiment Gramellini (2018), as well as the disagreement between $K^{+}-\text{Ca}$ interaction data and the GEANT4 interaction model sup .
The leading source of systematic uncertainty comes from the detector response (30.3%), with a major contribution from modeling ion recombination of electrons with argon. This analysis relies heavily on identifying the stopping $K^{+}$ and $\mu^{+}$ ($\pi^{+}$) signatures through their Bragg peak, which itself depends on the charge profile of the track in a highly ionizing regime. Because of this, the mismodeling of ion recombination, which alters the calorimetry in the detector, plays a particularly large role in the total error budget for the measurement. The modeling of neutrino interactions, the BNB neutrino flux, and particle reinteractions in the detector, including kaon inelastic interactions, lead to systematic uncertainties of 12.1%, 11.7%, and 10.4% respectively. The effect of statistics in the estimation of the selection efficiency (6.1%), the number of target nuclei $N_{\text{target}}$ (1%), simulation statistical uncertainty (7.4%), and POT exposure (2%) are the remaining sources of uncertainty. The overall magnitude of the uncertainty (statistical and systematic) on our measurement is 55.3%, where the major contribution is due to the statistical uncertainty, with a reported value of 41.3%. Table 1 presents the measured cross section of the $\nu_{\mu}+\text{Ar}\rightarrow\mu^{-}+K^{+}+X$ interactions and the predictions obtained from various neutrino event generators.
@@
MicroBooNE Data 7.93 $\pm$ 3.27 (stat.) $\pm$ 2.92 (syst.)
@@
In conclusion, this letter presents the first measurement of the $\nu_{\mu}+\text{Ar}\rightarrow\mu^{-}+K^{+}+X$ cross section on argon, using the MicroBooNE dataset collected from the BNB flux with a POT of $6.88\times 10^{20}$. The results reported here represent a significant step forward for future studies aiming to refine background estimates for proton decay searches, particularly within the framework of GUTs and SUSY. To advance this analysis in future LArTPC measurements, it is necessary to increase statistics, reduce large detector systematics (especially those related to the recombination of electrons with argon nuclei), refine reconstruction algorithms for short tracks, and improve the identification of kaon tracks that undergo reinteractions. Furthermore, the uncertainty related to kaon reinteractions on argon nuclei could be refined by using data from LArIAT Gramellini (2018), and the DUNE prototypes at CERN Abud et al. (2024). More statistics will come from analyzing MicroBooNE’s full dataset with a total of 1.2 ×$\times 10^{21}$ POT. The future DUNE near detector data could also provide more accurate measurements of CC neutrino-induced $K^{+}$ production on argon nuclei enabling improved comparisons among the neutrino generator models. These high-precision measurements will enhance our understanding of rare neutrino interactions and final state interaction modeling, and will improve the background predictions for future nucleon decay searches at DUNE, which currently rely on model predictions.

### v3

@@ Abstract
[…]
@@
Upcoming large long-baseline neutrino oscillation experiments, such as the Deep Underground Neutrino Experiment (DUNE) Abi et al. (2020) and Hyper-Kamiokande Abe et al. (2018), have the sensitivity to search for proton decay in the range of $10^{32}$ to $10^{39}$ years. Cherenkov detector experiments like Super-Kamiokande Abe et al. (2014, 2017, 2020) and Hyper-Kamiokande are particularly suited for $p\rightarrow e^{+}\pi^{0}$ decay searches. However, for water Cherenkov detectors, the $p\rightarrow\nu K^{+}$ decay channel needs to be indirectly detected since the momentum of the produced $K^{+}$ (339 MeV/c) is below the Cherenkov detection threshold in water (560 MeV/c). Furthermore, backgrounds from atmospheric neutrino interactions, such as the $\nu_{\mu}n\rightarrow\mu^{-}K^{+}n$ and $\nu p\rightarrow\nu K^{+}\Lambda$ channels, are challenging to distinguish in water Cherenkov detectors.
@@
The first measurements of charged-current (CC) and neutral-current (NC) kaon production, reported with the Argonne National Laboratory Barish et al. (1974) and GARGAMELLE bubble chambers Deden et al. (1975) in 1974 and 1975, observed $\mathcal{O}$(10) events. Later, in the 1980’s, the Brookhaven National Laboratory bubble chamber reported $\mathcal{O}$(10) CC and NC events with strange particle production in neutrino-hydrogen and neutrino-deuterium interactions Jensen (1964); Baker et al. (1981), and the Fermilab bubble chamber reported $\mathcal{O}$(10) events and a measured cross section for CC neutrino interactions $\nu_{\mu}+n\xrightarrow{}\mu^{-}+\Lambda+K^{+}$ on a deuterium target for energy ranges between 10 GeV and 250 GeV Son et al. (1983). Most recently, the MINERvA experiment measured $K^{+}$ production in NC and CC modes on a hydrocarbon target, reporting $\mathcal{O}$(100) and $\mathcal{O}$(1000) events respectively Marshall et al. (2016, 2017); Marshall (2016). In addition, MINERvA reported the first evidence of coherent kaon production Wang et al. (2016); Marshall (2016).
@@
As the first quantitative measurement of $K^{+}$ production on argon nuclei, this analysis will impact nuclear, neutrino, and proton decay research on two fronts. First, it provides background constraints in future $p\rightarrow\nu K^{+}$ nucleon decay experiments. Second, this measurement provides input to improve neutrino generator models. In the case of bound proton decay, the $K^{+}$ is generated within the nucleus and may interact with other nucleons as it escapes. These interactions can decrease the detected momentum of the $K^{+}$ in the detector and may also eject nucleons from the nucleus.
Charged-current neutrino-induced $K^{+}$ production occurs through three interaction modes: (1) single kaon production, where a CC interaction produces a single kaon with a production energy threshold of $E_{\nu_{\mu}}$ $\geq$ 0.79 GeV, and strangeness is not conserved ($|\Delta S|=1$); (2) associated kaon production, where both a kaon and a hyperon are produced, with a threshold of $E_{\nu_{\mu}}$ $\geq$ 1.2 GeV, and strangeness is conserved; and (3) coherent production, where only one kaon is produced in the final state, and the target nucleus remains intact after the neutrino interaction. Figure 1 shows an event candidate of CC $\nu_{\mu}$ kaon production recorded in MicroBooNE data. The kaon decays at rest into a 388 MeV energy muon, which subsequently decays into a Michel electron.
@@
The MicroBooNE detector Fleming (2012); Acciarri et al. (2017) consists of an 85 tonne LArTPC and a photon detection system comprised of 32 photomultiplier tubes (PMTs). The detector identifies neutrino interactions by detecting ionization and scintillation light produced by charged particles moving through the detector. Ionization charge is recorded across three wire planes (two induction, one collection) oriented at different angles (with induction planes at $\pm 60^{\circ}$ relative to the vertical collection plane), and 3 mm wire-spacing. This configuration allows the generation of mm-scale resolution, three-dimensional images of neutrino interactions. Scintillation light collected by the PMTs provides ns-scale timing information, essential for identifying neutrino interactions that are synchronized with the BNB, and for rejecting cosmic-ray backgrounds.
@@
In MicroBooNE, the LArSoft framework Snider and Petrillo (2017) provides tools for simulating the neutrino flux, neutrino interactions, particle propagation, and detector response including ionization and scintillation processes as well as reconstruction of neutrino events. The BNB neutrino flux through the MicroBooNE detector is simulated using the MiniBooNE flux model Aguilar-Arevalo et al. (2009) adapted for MicroBooNE’s location along the beamline. GENIE v3.0.6 (G18_10a_02_11a) Andreopoulos et al. (2010, 2021) is the event generator used to simulate neutrino-argon interactions, including both single and associated kaon production. In GENIE, associated kaons are produced via individual resonances simulated by the Rein-Sehgal model Rein and Sehgal (1981). Kaons produced via deep inelastic scattering are simulated by the Koba-Nielsen-Olesen (KNO) scaling parameterization Koba et al. (1972). For higher energies (hadronic invariant mass W >> 2.3 GeV), the AGKY model Yang et al. (2009) and PYTHIA6 Sjöstrand et al. (2006) are used. The hadronization models are tuned to match strange particle production ($\Lambda$ and $K^{0}_{S}$) observed in experiments such as the Big European Bubble Chamber Jones et al. (1993); Allasia et al. (1983); Willocq et al. (1992); Bosetti et al. (1982) and the Fermilab bubble chamber Baker et al. (1986); DeProspo et al. (1994). The single kaon production model Rafi Alam et al. (2010) is implemented in GENIE, generating events in the channels $\nu_{l}+n\rightarrow l^{-}+n+K^{+}$, $\nu_{l}+p\rightarrow l^{-}+p+K^{+}$, and $\nu_{l}+n\rightarrow l^{-}+p+K^{0}$. The region of validity assumed for the model is for $E_{\nu}\leq$ 2 GeV Rafi Alam et al. (2010); Alvarez-Ruso et al. (2013), while the GENIE authors note that the model can be applied to higher energies as well Alam et al. (2015).
@@
Particle trajectories and interactions within the detector are simulated using GEANT4 v4_10_3_p03c using the QGSP_BERT hadron physics list Agostinelli et al. (2003); Allison et al. (2006, 2016), which models particle propagation in the detector. Simulated neutrino interactions are overlayed with data events collected using an unbiased trigger that operates in anti-coincidence with the beam, enabling data-driven modeling of cosmic ray and detector noise. Both simulation and data reconstruction use the Pandora framework Acciarri et al. (2018), which uses a multi-algorithm approach to pattern recognition and event reconstruction in LArTPCs. Pandora creates clusters from wire hits, identifies the neutrino vertex, matches clusters across wire planes to form 3D objects, classifies reconstructed particles as tracks or showers, and establishes particle hierarchies by identifying parent-daughter relationships among particles.
@@
For the $K^{+}$ event selection, we implemented a boosted decision tree (BDT) classifier Hoecker et al. (2009). The BDT classifier inputs include reconstructed variables such as $\chi^{2}$ values which compare the expected $dE/dx$ for kaon ($\chi^{2}_{k}$), proton ($\chi^{2}_{p}$), pion ($\chi^{2}_{\pi}$), and muon ($\chi^{2}_{\mu}$) hypotheses with the observed data for each wire plane, the combined 3-wire-plane $\chi^{2}$, the daughter track length, and the log-likelihood ratio particle identification Abratenko et al. (2021). We trained the BDT classifier to identify true $K^{+}$ that decay to $\mu^{+}\nu_{\mu}$ or $\pi^{+}\pi^{0}$, using Monte Carlo (MC) simulation samples of reconstructed neutrino interactions that meet the preselection criteria described above. To train the BDT classifier, a background sample corresponding to $3.67\times 10^{20}$ POT (equivalent to 10% of the MicroBooNE BNB simulation) was used. For the simulation of the signal, a dedicated high-statistics sample of single and associated $K^{+}$ production was used. We optimized the BDT classifier score selection criteria by maximizing the product of efficiency and purity. Figure 2 shows the BDT classifier score distribution, with the orange arrow at 0.41 representing the selected BDT score cut. Events with a score higher than the BDT selection are kept as $K^{+}$ candidates, and 10 such $K^{+}$ candidates are identified in the data. The total number of background events predicted from neutrino-induced interactions, cosmic-rays recorded during neutrino-triggered beam-spills, and off-beam data is 2.21 events. The simulation predicts a selection efficiency of 3.95%, and an associated selection purity of 79%. The primary background category for $K^{+}$ candidates consists of protons generated at the neutrino vertex that interact with an argon nucleus, producing secondary particles at the proton endpoint that could mimic a $K^{+}$ decay product. This background category comprises 2.08 out of 2.21 predicted background events.
Figure 3 shows the reconstructed daughter track length distribution of the $K^{+}$ candidates selected in MicroBooNE data, including the event shown in Fig. 1. The stacked histogram shows the expected daughter track length distribution from the MicroBooNE BNB simulation. Due to the two-body decay nature of the $K^{+}$ branching ratios analyzed in this letter, the resulting monoenergetic $\mu^{+}$ ($\pi^{+}$) provides a powerful discriminant, as it leads to a fixed track length that can be measured with high accuracy and is not present in background events. The expected track length for $\mu^{+}$ ($\pi^{+}$) accumulates around 53 cm (30 cm). The $K^{+}$ candidates selected by the BDT selection in Fig. 3 (red triangles) exhibit a pile-up in the region compatible with $K^{+}$ decay.
@@
Figure 3: Reconstructed daughter track length distribution. The stacked histogram shows the MC simulation with its statistical uncertainty (gray band). The black dots represent the MicroBooNE data. The red triangles are the $K^{+}$ candidates selected in MicroBooNE data by the BDT, where three data points lie close to 60 cm, nearly overlapping.
To evaluate the accuracy of the background modeling, two sideband regions were defined using a 2-dimensional distribution of the BDT classifier score versus daughter track length. A far sideband region is defined to include tracks with low BDT classifier scores (BDT << 0) and track lengths less than 40 cm. A near sideband region selects tracks with BDT classifier scores close to the signal region (BDT << 0.41) and across all track length ranges, but excludes tracks in the far sideband region. The data/MC comparisons of the near and far sidebands show agreement for reconstructed kinematic variables such as $K^{+}$ candidate track length, $\phi$ (azimuthal angle around the beam direction), and $\theta$ (angle of the track with respect to the beam direction). For the $K^{+}$ candidate track length data/MC comparison, $\chi^{2}/\text{ndf}$ values (including statistical uncertainties) of 16.87/24 and 17.93/24 were obtained for the near and far sidebands. The agreement with the data suggests that the simulation is able to adequately model the background in the cross-section extraction. More information about the sideband comparisons is provided in the Supplemental Material sup .
This analysis presents the measurement of the single-bin flux-integrated cross section for $\nu_{\mu}+\text{Ar}\rightarrow\mu^{-}+K^{+}+X$ interactions, where muon neutrinos interact with Ar and produce one $\mu^{-}$, one $K^{+}$ and $X$ (any number of other hadrons in the final state) on argon nuclei. The cross section ($\sigma$) was found to be 7.93 $\pm$ 3.22 (stat.) $\pm$ 2.83 (syst.) $10^{-42}\text{cm}^{2}/{\rm nucleon}$, and it was calculated using the following expression:
@@
where $N$ represents the number of selected signal events (10), $B$ is the number of expected background events (2.21), $\varepsilon$ is the selection efficiency ($3.95\%$), $N_{\text{target}}$ is the number of target nucleons ($4.13\times 10^{31}$), and $\Phi_{\nu_{\mu}}$ stands for the integrated BNB $\nu_{\mu}$ flux ($6.02\times 10^{11}\nu/\text{cm}^{2}$). The selection efficiency accounts for all $K^{+}$ decays within the fiducial volume.
The systematic uncertainties in the measured cross section are evaluated by reweighting and generating simulated events to account for each source of uncertainty Abratenko et al. (2022a). For the detector-related uncertainties, a dedicated simulation is used to model the variation in detector responses Abratenko et al. (2022b); Calcutt et al. (2021). In contrast, uncertainties from neutrino interactions and BNB flux modeling are treated with a reweighting technique. To evaluate the uncertainty associated with kaon reinteractions within the detector, we implement a methodology similar to that presented by MINERvA Marshall et al. (2016). For each selected $K^{+}$ candidate track, a weight $W_{\text{inel}}$ is calculated using the following expression to estimate the kaon inelastic reinteraction uncertainty:
@@
where $\rho$ is the density of the material, $x$ is the track length, $\sigma^{\text{tot}}_{\text{geant}}$ and $\sigma^{\text{inel}}_{\text{geant}}$ are the $K^{+}$ total and inelastic GEANT4 cross sections, and $\sigma^{\text{tot}}_{\text{var}}$ and $\sigma^{\text{inel}}_{\text{var}}$ are the $K^{+}$ total and inelastic cross sections estimated by applying a variation of $\pm$40% to the $\sigma^{\text{inel}}_{\text{geant}}$. The 40% variation is motivated by observed disagreement between experimental measurements of $K^{+}-\text{Ca}$ interactions Friedman et al. (1997) and the predictions from the GEANT4 interaction model shown in the supplemental material sup .
The leading source of systematic uncertainty comes from the detector response (30.3%), with a major contribution from the modeling of ion recombination of electrons with argon. This analysis relies on identifying stopping $K^{+}$ and $\mu^{+}$ ($\pi^{+}$) through their Bragg peaks, which are characterized by a sharp rise in energy deposition near the end of the particle track. These effects occur in a highly ionizing regime, where ion recombination effects are most pronounced. As the charge profile is impacted by recombination, any mismodeling can distort the Bragg peak shape and affect particle identification. The modeling of neutrino interactions, the BNB neutrino flux, and particle reinteractions in the detector, including kaon inelastic interactions, lead to systematic uncertainties of 3.7%, 11.7%, and 10.5% respectively. The effect of statistics in the estimation of the selection efficiency (6.1%), the number of target nuclei $N_{\text{target}}$ (1%), MC statistical uncertainties on the background estimation (7.4%), and POT exposure (2%) are the remaining sources of uncertainty. The overall magnitude of the uncertainty (statistical and systematic) on our measurement is 54.1%, where the major contribution is due to the statistical uncertainty, with a reported value of 40.6%. Table 1 presents the measured cross section of the $\nu_{\mu}+\text{Ar}\rightarrow\mu^{-}+K^{+}+X$ interactions and the predictions obtained from various neutrino event generators with the NUISANCE framework Stowell et al. (2017).
@@
MicroBooNE Data 7.93 $\pm$ 3.22 (stat.) $\pm$ 2.83 (syst.)
@@
In conclusion, this letter presents the first measurement of the $\nu_{\mu}+\text{Ar}\rightarrow\mu^{-}+K^{+}+X$ cross section on argon, using the MicroBooNE dataset collected from the BNB flux. The results reported here represent a significant step forward for future studies aiming to refine background estimates for proton decay searches, particularly within the framework of GUTs and SUSY. To advance this analysis in future LArTPC measurements, it is necessary to increase statistics, reduce large detector systematics (especially those related to the recombination of electrons with argon nuclei), refine reconstruction algorithms for short tracks, and improve the identification of kaon tracks that undergo interactions. Experiments such as SBND Acciarri et al. (2025), ICARUS Amerio et al. (2004), and the DUNE near detector will allow precision measurements of neutrino-induced $K^{+}$ production on argon, including differential cross-section measurements for the first time ever. Furthermore, the uncertainty related to kaon reinteractions on argon nuclei could be refined by using data from LArIAT Gramellini (2018), and the DUNE prototypes at CERN Abud et al. (2024). More statistics will come from analyzing MicroBooNE’s full dataset with a total of 1.2 $\times 10^{21}$ POT. The future DUNE near detector data could also provide more accurate measurements of CC neutrino-induced $K^{+}$ production on argon nuclei enabling improved comparisons among the neutrino generator models. These high-precision measurements will enhance our understanding of rare neutrino interactions and final state interaction modeling, and will improve the background predictions for future nucleon decay searches at DUNE, which currently rely on model predictions.